Find 3 consective odd integers such that 8 more than twice the first integer is equal to 19 less than three times the third integer.
Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

8 + 2(x-2) = 3(x+2) - 19
2x + 4 = 3x - 13
17 = x

The integers are 15, 17, and 19.